# In situ measurement of slow solar wind emerging from a pseudostreamer: a conjunction study with Parker Solar Probe and Solar Orbiter

### v2

@@ 3.2 Magnetohydrodynamic Models
In order to directly compare the measured observables to the model, we sample the MHD modeled observables of interest at PSP’s location in 3D for the full trajectory (Figure 3) and compare with the radial magnetic field, density, and velocity measurements from PSP which are taken fully within the coronal model regime. A full Carrington map of the radial cuts of the magnetic field, density, and velocity observables compared with in-situ measurements can be found in the Appendix (see Section 9 Figure 9). We find very good agreement between the measured and modeled observables, validating that our model is properly functioning and is usable to trace magnetic field lines and estimate photospheric footpoints. In Figure 3, we show the correlation between the PFSS and MHD estimated footpoints: 0.998 in longitude and 0.617 in latitude for our chosen model with a source surface height of 2.0. The MHD tracing has no assumptions of ballistic propagation or source surface height and thus provides a strong comparison point with our estimated PFSS footpoints. This, coupled with the strength of the MHD model in reproducing measured plasma parameters, allows for confidence in the footpoints as estimated via the PFSS and MHD models.
@@ 4 Observations and Discussion
We combine data from both PSP and Solar Orbiter during PSP E11 with modeling results to understand the origins of the SSW and SASW. In Figure 2, we see that both spacecraft had sharp HCS crossings that map very close to each other (Figure 1) while traveling in opposite directions in the solar corotating frame: PSP on February 25, 2022 and Solar Orbiter on March 1, 2022. The HCS crossing is sharper in the PSP measurements as PSP moves at a much larger tangential speed than Solar Orbiter.
@@
We use ballistic mapping to map the observed plasma back to the source surface and plot the data as a function of Carrington longitude rather than as a timeseries. This allows us to more easily directly compare structures observed by both spacecraft and relate SWA/HIS composition measurements to PSP in situ signatures. In this section, we outline the observations and discuss their implications in determining the source regions of the solar wind.
4.1 Heavy Ion and Elemental Composition Signatures
In situ elemental composition measurements can be used to connect heliospheric structures to their solar origin by comparing them to characteristic elemental abundances of coronal sources. Similarly, ion composition reflects characteristics of the source region’s thermal structure, as well as the properties of its thermodynamic evolution. By combining the elemental composition and charge state measurements from Solar Orbiter (measurements PSP does not provide) with PSP measurements close to the Sun, which degrade with radial distance (such as Alfvénicity), we can more effectively relate heliospheric structure to their coronal sources.
As seen in Figure 4, the ion (, ) and elemental () composition measurements vary greatly over the time period of interest. The ratio in panel (a) is scaled to its photospheric value according to (()/) where = 0.0589 (Asplund et al., 2021). From this point on, discussion of the ratio will refer to the normalized (()/) value. The normalized ratio has been shown to relate to coronal source region properties and alongside modeling is a good tracer of solar wind source. The yellow-shaded region shows ratios between 1 and 2 (photosphere-like composition typical of CH regions) while the grey region covers ratios from 2 to 4 (coronal-like composition typical of non-CH wind). Similarly, the ion charge state ratios in panel (b) are a diagnostic of a coronal temperature, allowing us to probe the thermal properties of the stream’s coronal source. We include a dashed line in panel (b) at 0.145 which is the cutoff for the ratio from Wang (2016) between streamer (high ) and CH wind.
We see variance in the ratio, which implies a combination of source regions for the plasma that is measured throughout the encounter. We primarily see ratios ranging between 2 and 4, which is usually the case for active region outflow as found in Feldman & Laming (2000); Widing & Feldman (2001). Similar to the ratio, ion charge state ratios provide a key tracer of the plasma the spacecraft is measuring, indicating the electron temperature and density along the path the plasma parcel took through the corona. The majority of this time period consists of low and ratios, which imply a low characteristic electron source temperature. Throughout the encounter, the ratio varies between 0.022 and 0.44, while the ratio shows a far wider variance between 0.42 and 14.9. The ratio ranges between 0.80 and 5.25 for this time period.
In panels (c), (d), and (e) of Figure 4 and in Table 1, we investigate the correlations between the charge state and elemental abundance ratios. The data points are colorized according to the time interval of interest shown in the left-hand panel. The ion ratios ( and ) are strongly correlated with each other, with a Spearman correlation coefficient of 0.78, but not with the ratio (Figure 4). Both the and ratios are indicators of electron temperature with some radial evolution effects and as such, are expected to follow the same trends as they have similar freeze-in heights (Zhao et al., 2017). The and ratios have Spearman correlation coefficients of 0.15 and 0.33 with , respectively. The lack of overall correlation between the charge state ratios and elemental composition implies the presence of a variety of solar wind sources. There has not been a strong correlation observed between and solar wind speed as seen by both SWA/HIS and ACE/SWICS (Livi et al., 2023) and so we would not necessarily expect a strong correlation with the charge state ratios and . This could be because the characteristic source region temperature can be transient in nature (heating and cooling occurring at noticeable timescales) while the FIP effect takes place on a longer timescale, or the outflow effects are important – the density, temperature, and bulk speed that set the ion ratios during outflow varies across source regions with similar .
When looking at the correlations for the different time periods of interest, we see some distinct groupings. The FSW typically shows far lower ion ratios than the other types of wind, especially when compared to the SSW and HCS wind. It shows a strong correlation between and but a low negative correlation between the ion ratios and . The SSW shows some correlation between the ion ratios, almost no correlation between and and a low negative correlation for and . Similar to the FSW, the SASW has a higher correlation between the charge state ratios. Looking at the correlation plots, we see two groups of SASW, most distinct in panel (c). In all three panels, there is a lower left blob of points more similar to FSW-like characteristics and an upper right blob of points found almost between the HCS and SSW blobs. The SSW also shows some grouping that is found mainly at the inner boundary between the SASW blobs with some points closer to the HCS plasma, while the HCS grouping is on the outer right edge of the SASW group. Similar to Zhao et al. (2017); Rivera et al. (2021), we see some of this characteristic ‘Outlier’ SSW in panel (c) of Figure 4 where the ratio is much lower than expected in comparison to the ratio. This is slow wind thought to have streamer or active region origins and is strongly correlated with the solar cycle (Rivera et al., 2021).
| Wind Type | vs. | vs. | vs. |
| All | 0.78 | 0.33 | 0.15 |
| HCS | 0.43 | -0.38 | -0.45 |
| FSW | 0.85 | -0.27 | -0.19 |
| SSW | 0.50 | 0.012 | -0.27 |
| SASW | 0.66 | 0.19 | -0.03 |
In the FSW, we see low and ratios during the region, and a dip in these ratios just before and after the regions showing a transition to cooler, less dense coronal hole plasma. The ratio during this region has photosphere-like abundance levels with the sharpest dip in the center of the region – indicative of CH type plasma.
The ratio in the SSW (green) region is well above the threshold of 0.145, implying hot coronal conditions indicative of non-CH origins. There is also a spike in the ratio at the boundary between the SSW and SASW region implying a pseudostreamer crossing. These composition measurements support our claim that this SSW wind period originates from a pseudostreamer due to the spike in at the boundary between the Alfvénic and non-Alfvénic SSW and the high ion charge state ratios in this region.
The SASW region has a huge variance in the ratio. The ratio increases and decreases through the center of the region indicating some variation in the structure the spacecraft is connected to. Both ion charge state ratios show large variation in this region from more CH-type wind to streamer-like wind. As discussed previously, there seem to be two groupings of composition results for the SASW. This is due to the overlapping in the backmapping in this region, as seen in Figure 1.
4.2 Particle Measurements
In Figure 5 we show the PSP particle measurements. In panel (a), we show the alpha abundance ratio () and in panel (b) we show the differential velocity normalized by the Alfvén speed (). In both panels, the color corresponds to the solar wind bulk velocity and the black line shows the data binned by 1° in longitude. We also compare the scaled proton and alpha abundance as a function of wind type in panel (c), using the dashed lines to separate between the high () and low () abundance regimes as defined by Kasper et al. (2007, 2012). As in all Figures, we shade the regions of interest and discuss interesting observations below.
In Figure 5, we see that overall, the abundance ratio during E11 rarely crosses (and never in the binned data) above the typical ‘high abundance’ threshold of 0.045 (Kasper et al., 2007, 2012). There is some variance in abundance throughout the encounter, however the ratio remains low for the majority of this period. There is a strong correlation between the proton () and alpha particle () bulk velocity with a Spearman correlation of 0.9. The correlation between the differential velocity and the solar wind velocity is weak (correlation of 0.50) and is much stronger between the fast solar wind and differential velocity (0.51) than for the slow wind (0.27). We find that for the entire time frame, the alpha-to-proton differential velocity is below the Alfvén speed, (panel (b)). This is due to alpha-proton instabilities driving wave-particle interactions to hit the instability threshold, leading to an upper bound on the differential speed of (Gary et al., 2000).
At the HCS crossing, we see an increase in proton density in both the PSP and Solar Orbiter data (Figure 2) along with a dip in helium abundance in panel (b) of Figure 5. There are two clusters of abundance ratios in panel (c): one just above the low abundance threshold and one just below. There is also a wide range of scatter in the differential velocity measurement, with the binned value around 0.25.
In the FSW region, we see both intermediate and very low helium abundances along with differential velocity values ranging between 0.25 and 0.6, and binned values that are higher than in all other regions of interest. In panel (c), we see that the fast wind points fall in the lower left of the plot, similar to the composition data seen in Figure 4. Typically, we expect to see an enhancement in helium abundance in a fast wind regime, which is not the case in the PSP data shown here.
In the SSW region, we see the lowest velocities as measured by PSP (187 ) during this time range and also a sharp dip in the normalized velocity differential with a drop of 97.7%. This drop also corresponds to the lowest helium abundance ratio in the binned data measured throughout this interval (0.004). Low helium abundance wind is characteristic of slow wind from streamers (Kasper et al., 2007), another important piece of evidence for this slow solar wind region to be from a pseudostreamer. In panel (c), we see two groupings of SSW plasma: one with low and one with high proton and alpha densities, both of which are below the low abundance threshold.
The SASW consists of both alpha particle intermediate and poor plasma populations implying that it likely originates from multiple source regions. We can see these two groupings in panel (c) where one group shows more HCS-like abundance ratios, and the other shows more FSW-type ratios.
@@ 4.3 Modeling
In Figure 6, we see the first indication that the SSW period came from a streamer crossing – the jump in photospheric footpoints, which indicates the spacecraft crossed a region of closed field connecting two CHs. The background and observed polarity are unipolar, which indicates this is a pseudostreamer rather than a helmet streamer (crossing between two different polarity CHs). In Figure 7, we investigate the connectivity of the SSW stream by taking latitude (panel (c)) and longitude (panel (d)) slices through the 3D MHD solution. The latitude and longitude chosen are shown by the horizontal and vertical dashed lines in Figure 6 respectively. Taking these slices through the model, allows for the determination of the magnetic field configuration around the footpoints of the region, and when used in conjunction with the low expansion factor and slow wind speed indicates that the spacecraft flew through a pseudostreamer during this period (Riley et al., 2015; Wallace et al., 2020). Panel (c) of Figure 7 shows the jump in footpoints (green diamonds) across a closed field region, indicating that the footpoints are connected to the edges of a unipolar streamer. The trajectory of PSP at the source surface (2.0 ) is shown with diamonds that are colored by the measured proton density. We see a narrow density enhancement corresponding to the jump in footpoints in the Carrington frame as PSP crosses the unipolar streamer in both the modeled magnetic field, and in situ observation. In panel (d), we show a longitude slice through the MHD model, where two footpoints (green diamonds) are connected to the edge of a unipolar streamer: one footpoint in front of and one behind the longitude slice. This is because PSP is flying through the plane of the sky during this period. Combining the modeling results, with the in situ observations of a narrow density enhancement, drop in speed, uniform polarity and low expansion factor indicate that this SSW stream came from PSP’s crossing of a pseudostreamer structure.
@@
The SASW stream shows some similar properties to the SSW stream and some to the FSW stream. This implies a combination of source regions – the periphery of the pseudostreamer connecting into a positive polarity CH boundary. In panel (b) of Figure 7, we show the SASW stream emerging from the boundary of an equatorial CH. This period has an elevated expansion factor (panel (c) of Figure 6), which is characteristic of wind streaming from the over-expanded pseudostreamer/CH boundary (Panasenco & Velli, 2013; Panasenco et al., 2013; Sheeley et al., 2013; Wang, 2013; Panasenco et al., 2019).
@@ 5 Results
By using a combination of methods, we are able to probe the source regions of the solar wind by connecting models to in situ data. The connection between modeling and in situ data provides insight from both a theoretical and observational point of view as to the origins of the solar wind. While the modeling methods do have user-defined inputs, when combined with in situ observations, they provide strong evidence for source regions of the slow solar wind.
@@
As expected from von Steiger et al. (2000); McComas et al. (1998, 2008), we find that the fast wind stream maps to a relatively large coronal hole region, with low ion charge state ratios and photospheric-like Fe/O abundances, characteristic of coronal hole wind.
@@
The slow wind has been shown to have large variation in plasma properties and be associated with the HCS (Antiochos et al., 2011), S-Web (Chitta et al., 2023; Lynch et al., 2023), and pseudostreamers (Wang et al., 2012; Riley & Luhmann, 2012). Similar to Wang et al. (2012); Riley & Luhmann (2012), our combination of modeling and in situ measurements show that the SSW region of interest originates from the stalk of a pseudostreamer. We see an enhancement in ion charge state ratios, FIP bias, and proton density around this region, which is typical of streamer wind.
The high ratio during this period is indicative of the FIP effect which has been shown to require plasma to be on closed loops and remain in the corona for longer time periods to fractionate (Geiss et al., 1995; von Steiger et al., 2000; Laming, 2015). This time period has a spike in proton density, low Alfvénicity in both PSP and Solar Orbiter observations, the lowest helium abundance of the entire time period and a sharp drop in the normalized velocity differential. We also see high ion charge state ratios which implies a high electron temperature of the source region.
The spike in charge state ratios at the boundary between the SSW and SASW regions coupled with the jump in footpoints seen in Figure 6 and density enhancement in both PSP and Solar Orbiter data (Figure 2) without crossing the HCS is indicative that this non-Alfvénic SSW plasma originated from a pseudostreamer stalk.
4.
The slow Alfvénic wind during this time period is shown to have two populations of plasma with different characteristics: one that shows more FSW-type characteristics and the other showing more streamer-like properties. Similar to results from D’Amicis et al. (2021), the SASW is thought to originate from an over-expanded pseudostreamer connected to a CH boundary.
The SASW shows coronal-like abundance ratios with a spike in the center of the region, showing some variance and perhaps evolution in the structure the spacecraft are magnetically connected to. The ion abundance ratios bound this region with a large range of measured values. The correlation plots show two groupings: one with high charge state ratios and low FIP enhancement and the other with low charge state ratio and photospheric level elemental composition. This regime also shows variance in alpha-to-proton abundance with a grouping similar to the HCS abundance and one similar to the FSW abundance. These groupings along with the modeling results in Figure 6 imply the SASW comes from a pseudostreamer periphery that connects into a positive polarity coronal hole, and the CH boundary itself. Over expansion can occur due to the presence of coronal pseudostreamers, which when coupled with low coronal non-monotonic expansion can slow the solar wind leading to similar properties to the FSW, but much slower wind speeds (Panasenco & Velli, 2013; Panasenco et al., 2019; D’Amicis et al., 2021). It is important to note that the neighboring SSW and SASW are difficult to disambiguate in the Solar Orbiter data due to the velocity gradient and ballistic mapping as Solar Orbiter is further from the Sun leading to overlap in the backmapping of these two time periods.

### v3

@@ 3.2 Magnetohydrodynamic Models
In order to directly compare the measured observables to the model, we sample the MHD modeled observables of interest at PSP’s location in 3D for the full trajectory (Figure 3) and compare with the radial magnetic field, density, and velocity measurements from PSP which are taken fully within the coronal model regime. A full Carrington map of the radial cuts of the magnetic field, density, and velocity observables compared with in-situ measurements can be found in the Appendix (see Section 9 Figure 9). We find very good agreement between the measured and modeled observables, validating that our model is properly functioning and is usable to trace magnetic field lines and estimate photospheric footpoints. In Figure 3, we show the correlation between the PFSS and MHD estimated footpoints: 0.998 in longitude and 0.674 in latitude for our chosen model with a source surface height of 2.0. The MHD tracing has no assumptions of ballistic propagation or source surface height and thus provides a strong comparison point with our estimated PFSS footpoints. This, coupled with the strength of the MHD model in reproducing measured plasma parameters, allows for confidence in the footpoints as estimated via the PFSS and MHD models.
@@ 4 Observations and Discussion
We combine data from both PSP and Solar Orbiter during PSP E11 with modeling results to characterize the plasma emerging from FSW and SASW. In Figure 2, we see that both spacecraft had sharp HCS crossings that map very closely to each other (Figure 1) while traveling in opposite directions in the solar corotating frame: PSP on February 25, 2022 and Solar Orbiter on March 1, 2022. The HCS crossing is sharper in the PSP measurements as PSP moves at a much larger tangential speed than Solar Orbiter.
@@
We use ballistic mapping to map the observed plasma back to the source surface and plot the data as a function of Carrington longitude rather than as a timeseries. This allows us to more easily directly compare structures observed by both spacecraft and relate SWA/HIS composition measurements to PSP in situ signatures. In this section, we outline the source regions of the streams of interest and then describe the associated in situ characteristics.
@@ 4.1 Modeling
The second FSW stream (green footpoints) emerges from a CH region near 125° in longitude. Due to the proximity of this FSW stream to the HCS in the model, the expansion factor and footpoint field strength are much larger than for the other FSW (pink) region. In Figure 5, we see these footpoints lie within a dark region in SDO/AIA EUV observations, which combined with the low charge state and FIP bias of this period, indicates CH wind.
@@
The SASW stream shows some similar properties to the FSW stream. In panel (c) of Figure 5, we show the SASW stream emerging from the boundary of an equatorial CH. This period has an elevated expansion factor (panel (c) of Figure 4), which is characteristic of wind streaming from the over-expanded CH boundary (Panasenco & Velli, 2013; Panasenco et al., 2013; Sheeley et al., 2013; Wang, 2013; Panasenco et al., 2019).
4.2 Heavy Ion and Elemental Composition Signatures
By combining the elemental composition and charge state measurements from Solar Orbiter (measurements PSP does not provide) with PSP measurements close to the Sun, which degrade with radial distance (such as Alfvénicity), we can more effectively relate heliospheric structure to their coronal sources.
In Figure 6, we outline the ion (, ) and elemental () composition measurements. The ratio is scaled to its photospheric value according to (()/) where = 0.0589 (Asplund et al., 2021). From this point on, discussion of the ratio will refer to the normalized (()/) value. The normalized ratio has been shown to relate to coronal source region properties and alongside modeling is a good tracer of solar wind source. Ratios between 1 and 2 are indicative of photosphere-like composition typical of CH regions, while ratios from 2 to 4 show coronal-like composition typical of non-CH wind. Similarly, the ion charge state ratios in panels (a) and (b) are a diagnostic of a coronal temperature, allowing us to probe the thermal properties of the stream’s coronal source. We include a dashed line in panel (a) at 0.145 which is the cutoff for the ratio from Wang (2016) between streamer (high ) and CH wind.
We see variance in the ratio for all wind types, which implies a combination of source regions for the plasma that is measured throughout the encounter. We primarily see ratios ranging between 2 and 4, which is usually the case for active region outflow as found in Feldman & Laming (2000); Widing & Feldman (2001). The majority of this time period consists of low and ratios, which imply a low characteristic electron source temperature. Throughout the encounter, the ratio varies between 0.022 and 0.44, while the ratio shows a far wider variance between 0.42 and 14.9. The ratio ranges between 0.80 and 5.25 for this time period.
In panels (d), (e), and (f) of Figure 6 and in Table 1, we investigate the correlations between the charge state and elemental abundance ratios. The data points are colorized according to the time interval of interest shown in the left-hand panel. The ion ratios ( and ) are strongly correlated with each other, with a Spearman correlation coefficient of 0.75, but not with the ratio (Figure 6). Both the and ratios are indicators of electron temperature with some radial evolution effects and as such, are expected to follow the same trends as they have similar freeze-in heights (Zhao et al., 2017). The and ratios have Spearman correlation coefficients of 0.15 and 0.40 with , respectively. The lack of overall correlation between the charge state ratios and elemental composition implies the presence of a variety of solar wind sources. There has not been a strong correlation observed between and solar wind speed as seen by both SWA/HIS and ACE/SWICS (Livi et al., 2023) and so we would not necessarily expect a strong correlation with the charge state ratios and . This could be because the characteristic source region temperature can be transient in nature (heating and cooling occurring at noticeable timescales) while the FIP effect takes place on a longer timescale, or the outflow effects are important – the density, temperature, and bulk speed that set the ion ratios during outflow varies across source regions with similar .
When looking at the correlations for the different time periods of interest, we see some distinct groupings. The FSW streams typically shows far lower ion ratios than the other types of wind which is expected for CH wind. It shows a strong correlation between and but is relatively uncorrelated in the other panels. Similar to the FSW, the SASW has a higher correlation between the charge state ratios. Looking at the correlation plots, we see two groups of SASW, most distinct in panel (d). In all three panels, there is a lower left blob of points more similar to FSW-like characteristics and an upper right blob of points found nearer to the HCS wind.
Wind Type | vs. | vs. | vs. |
All | 0.75 | 0.40 | 0.15 |
HCS | 0.05 | -0.24 | -0.40 |
FSW (1) | 0.54 | 0.02 | -0.35 |
FSW (2) | 0.84 | 0.038 | -0.006 |
SASW | 0.71 | 0.16 | -0.08 |
We see low and ratios during the FSW streams, and a dip in these ratios just before and after the regions showing a transition to cooler, less dense coronal hole plasma. The ratio (see Figure 2) during this both FSW streams has photosphere-like abundance levels with the sharpest dip in the center of the region – indicative of CH type plasma. For the charge state ratios and ratio, the FSW streams show incredibly low values indicating that the spacecraft were sampling wind from deep within a CH.
The SASW region has a huge variance in all ratios. The (see Figure 2) ratio increases and decreases through the center of the region indicating some variation in the structure the spacecraft is connected to. Both ion charge state ratios show large variation in this region from more CH-type wind to streamer-like wind. While the values for the SASW are not as distinctively low as for the FSW streams, they are far below the HCS wind indicating their origin from a cooler source region. As discussed previously, there seem to be two groupings of composition results for the SASW. This is due to the overlapping in the backmapping in this region, as seen in Figure 1.
4.3 Particle Measurements
In Figure 7 we show the PSP particle measurements. In panel (a), we show the scaled proton and alpha abundance as a function of wind type and in panel (b) the differential velocity normalized by the Alfvén speed () as a function of wind speed. The dashed lines in panel (a) separate between the high () and low () abundance regimes as defined by Kasper et al. (2007, 2012).
In Figure 7, we see that overall, the abundance ratio during E11 rarely crosses above the typical ‘high abundance’ threshold of 0.045 (Kasper et al., 2007, 2012). There is some variance in abundance throughout the encounter, however the ratio remains low for the majority of this period. There is a strong correlation between the proton () and alpha particle () bulk velocity with a Spearman correlation of 0.9. The correlation between the differential velocity and the solar wind velocity is weaker (correlation of 0.75) and is stronger between the fast solar wind and differential velocity (0.88 and 0.63 for FSW (1) and FSW (2)) than for the SASW (0.59). We find that the alpha-to-proton differential velocity is typically below the Alfvén speed, (panel (b)). This is due to alpha-proton instabilities driving wave-particle interactions to hit the instability threshold, leading to an upper bound on the differential speed of (Gary et al., 2000).
At the HCS crossing, we see an increase in proton density in both the PSP and Solar Orbiter data (Figure 2) along with a decreased in helium abundance in panel (a) of Figure 7. There are two clusters of abundance ratios: one just above the low abundance threshold and one just below. There is also a wide range of scatter in the differential velocity measurement, with the binned value around 0.25.
In the FSW regions, we see intermediate but primarily very low helium abundances. This is atypical for FSW but is seen in both streams discussed during this study and thus warrants further investigation. The differential velocity correlates well with wind speed as is expected (Mostafavi et al., 2022).
The SASW consists of both alpha particle intermediate and poor plasma populations implying that it likely originates from multiple source regions. We can see these two groupings in panel (a) where one group shows more HCS-like abundance ratios, and the other shows more FSW-type ratios. It shows differential streaming speeds very similar to the FSW populations and a strong correlation between and wind speed.
@@ 5 Results
By connecting models to in situ data, we are able to probe the characteristics of the solar wind emerging from CHs. The connection between modeling and in situ data provides insight from both a theoretical and observational point of view as to the origins of the solar wind. While the modeling methods do have user-defined inputs, when combined with in situ observations, they provide strong evidence to characterize the compositional signature of varying types of CH wind.
@@
As expected from von Steiger et al. (2000); McComas et al. (1998, 2008), we find that two fast wind (FSW) streams that maps to relatively large coronal hole regions, with low ion charge state ratios and photospheric-like Fe/O abundances, characteristic of coronal hole wind.
@@
The slow Alfvénic wind (SASW) during this time period is shown to have two populations of plasma with different characteristics: one that shows more FSW-type characteristics and the other showing more streamer-like properties. Similar to results from D’Amicis et al. (2021), the SASW is thought to originate from the over-expanded edges of a CH boundary.
The SASW shows coronal-like abundance ratios with a spike in the center of the region, showing some variance and perhaps evolution in the structure the spacecraft are magnetically connected to. The ion abundance ratios bound this region with a large range of measured values. The correlation plots show two groupings: one with high charge state ratios and low FIP enhancement and the other with low charge state ratio and photospheric level elemental composition. This regime also shows variance in alpha-to-proton abundance with a grouping similar to the HCS abundance and one similar to the FSW abundance. These groupings along with the modeling results in Figure 4 imply the SASW comes from the over-expanded boundary of a positive polarity coronal hole. Over expansion can occur due to the presence of coronal pseudostreamers, which when coupled with low coronal non-monotonic expansion can slow the solar wind leading to similar properties to the FSW, but much slower wind speeds (Panasenco & Velli, 2013; Panasenco et al., 2019; D’Amicis et al., 2021).